What is the molarity of a sodium hydroxide solution if 4.50 mL of the solution is titrated to the end


point, with 5.00 mL of acetic acid with 0.0830M?

1
Expert's answer
2022-03-10T10:45:00-0500

NaOH + HAc = NaAc + H2O;

C1 * V1 = C2 * V2;

C(NaOH) * V(NaOH) = C(HAc) * V(HAc);

C(NaOH) = C(HAc) * V(HAc) / V(NaOH) = 0.0830 * 5 / 4.5 = 0.0922 M.

Answer: 0.0922 M.